OpenStudy (anonymous):

solve by elimination or substitution: x^2-y^2=15 and x+y=1

OpenStudy (anonymous):

x^2-y^2=(x+y)(x-y)=15 and we have x+y=1 so 1x15=15 so x-y=15 (x-y)+(x+y)=2x=15+1 so x=8 y=-7